OpenStudy (anonymous):

i will medal and fan you! 1.(–10x3 + 30x – 20) ÷ (–5x + 5) 2.(8m7 – 10m5) ÷ 2m3 3.(–6m9 – 6m8 – 16m6) ÷ (2m3) these kinds of questions confuse the heck out of me please help and explain(:

OpenStudy (ajprincess):

is ur first question? .(–10x3 + 30x – 20) ÷ (–5x + 5) or else .(–10x2 + 30x – 20) ÷ (–5x + 5)?

OpenStudy (ajprincess):

@FullHouseSquidy

OpenStudy (anonymous):

these are the questions. sorry i wrote them down wrong: (-10x^3+30x-20) / -5x+5 (8m^7 – 10m^5) ÷ 2m^3 (–6m^9 – 6m^8 – 16m^6) ÷ (2m^3)

OpenStudy (ajprincess):

Let us try to do the first question. Can u factorise (-10x^3+30x-20)?

OpenStudy (anonymous):

no you cant

OpenStudy (ajprincess):

\(-10x^3+30x-20=-10x^3+3*10x-2*10\) do u see anything common to all three terms?

OpenStudy (anonymous):

10

OpenStudy (ajprincess):

right:) when u factor it out what remains?

OpenStudy (anonymous):

-2^3+3-2?

OpenStudy (anonymous):

oh wait its -z^3+3-2

OpenStudy (ajprincess):

\(-10x^3+3*10x-2*10=10(-x^3+3x-2)\) getting it?

OpenStudy (anonymous):

yeah i get that

OpenStudy (ajprincess):

good:) Can u factorise -5x+5?

OpenStudy (anonymous):

-5(x+1)

OpenStudy (ajprincess):

-5x+5=-5(x-1)

OpenStudy (anonymous):

oh ok

OpenStudy (ajprincess):

\[\frac{ -10x^3+30x-20 }{ -5x+5 }=\frac{ -10(x^3-3x+2) }{ -5(x-1) }=\frac{ -5*2( x^3-3x+2) }{ -5(x-1) }\] do u see anything common in both denominator and numerator?

OpenStudy (anonymous):

the -5 and x?